# On Benchmarking Code LLMs for Android Malware Analysis

### v1

@@ Abstract.
Large Language Models (LLMs) have demonstrated strong capabilities in various code intelligence tasks. However, their effectiveness for Android malware analysis remains underexplored. Decompiled Android code poses unique challenges for analysis, primarily due to its large volume of functions and the frequent absence of meaningful function names. This paper presents Cama, a benchmarking framework designed to systematically evaluate the effectiveness of Code LLMs in Android malware analysis tasks. Cama specifies structured model outputs (comprising function summaries, refined function names, and maliciousness scores) to support key malware analysis tasks, including malicious function identification and malware purpose summarization. Built on these, it integrates three domain-specific evaluation metrics—consistency, fidelity, and semantic relevance—enabling rigorous stability and effectiveness assessment and cross-model comparison. We construct a benchmark dataset consisting of $118$ Android malware samples, encompassing over $7.5$ million distinct functions, and use Cama to evaluate $4$ popular open-source models. Our experiments provide insights into how Code LLMs interpret decompiled code and quantify the sensitivity to function renaming, highlighting both the potential and current limitations of Code LLMs in malware analysis tasks.
doi: conference: ; June 25–28, 2025; Trondheim, Norwaybooktitle: Workshop on Large Language Model Supply Chain Analysis (LLMSC ’25) Co-located with ISSTA’25, June 25–28, 2025, Trondheim, Norway
@@ 1. Introduction
Recent advancements in Large Language Models (LLMs), driven by transformer-based architectures and self-supervised learning on massive corpora (Radford et al., 2018; Raffel et al., 2020; Devlin et al., 2019), have significantly improved natural language processing (NLP) tasks. Extending these successes to the programming domain, Code LLMs (i.e., specialized LLMs trained on large-scale code repositories) have emerged to effectively capture code syntax and patterns. Models such as CodeLlama (Roziere et al., 2023) and StarCoder (Li et al., 2023a) have demonstrated strong performance across software engineering tasks including code generation, code summarization, and code repair (Li et al., 2023b; Gu, 2023; Tian et al., 2024; Fang et al., 2024).
Despite these promising developments, applying Code LLMs to Android malware analysis remains challenging. First, the inherent complexity of decompiled Android code—including obfuscated function names, missing type information, and incomplete control structures (Meng and Miller, 2016; Shang et al., 2024)—significantly hinders the effectiveness of current models in code summarization. Such decompiled code diverges substantially from the clean and structured source code used in pre-training Code LLMs (Zheng et al., 2023). Second, accurate semantic interpretation remains difficult due to the high-level abstraction and diverse malicious behaviors present in Android malware (Wang et al., 2022). Lastly, the scarcity of reliable ground-truth labels at the function level complicates rigorous model evaluation (He et al., 2023b; Kantchelian et al., 2015; Salem, 2021). These challenges underscore the need for a structured evaluation framework that systematically assesses and compares Code LLM performance in real-world malware analysis scenarios.
@@
For the first question, we analyze the quality of LLM-generated outputs using our structured output format and domain-specific metrics. For the second, we systematically rename functions in the original decompiled code with the LLM-suggested names and measure the subsequent impact on evaluation metrics. Our findings reveal that while Code LLMs can generate informative function summaries, their understanding of maliciousness remains limited, highlighting substantial room for improvement. Among existing Code LLMs, instruction-tuned GPT-style models significantly outperform Seq2Seq models across different metrics; function renaming further enhances fidelity and consistency but may reduce semantic clarity, indicating a trade-off that warrants careful consideration.
@@ 2.1. Code Large Language Models
2.2. Learning-based Malware Analysis
LLM-powered Analysis. Recent efforts have explored LLM-powered malware detection, primarily operating within the established pipeline of conventional classifiers and leveraging LLMs in two ways: 1) querying LLMs with original features to generate detection outputs (Li et al., 2024), and 2) using LLMs to encode text-based semantic representations that enrich traditional feature spaces (Zhao et al., 2025). More advanced approaches leverage GPT-4o-mini’s code summarization capabilities for multi-level malware analysis (Walton et al., 2025), and improving malware detection via program slicing techniques and multi-tiered code reasoning for factual checking (Qian et al., 2025). Despite these advancements, the effectiveness of Code LLMs in fine-grained analysis remains uncharted, largely due to the lack of ground truth. Our work is orthogonal to existing studies, systematically benchmarking open-source Code LLMs in structured malware analysis tasks.
@@ 3.3.1. Consistency-based Metric
(1) $\text{JSD}(M^{\prime}_{\text{raw}},M^{\prime}_{\text{des}})=\frac{1}{2}D_{\text{KL}}(M_{\text{raw}}||M_{\text{avg}})+\frac{1}{2}D_{\text{KL}}(M_{\text{des}}||M_{\text{avg}})\,,$
@@ 3.3.3. Semantic-based Metric
We compare $A_{\text{LLM}}$ against the reference malware description $A_{\text{GT}}$ using three widely used text similarity metrics:
  1. a)
    BLEU (Papineni et al., 2002): Measures n-gram precision between $A_{\text{LLM}}$ and $A_{\text{GT}}$:
@@
    where $p_{n}$ is n-gram precision and $w_{n}$ are weighting factors.
  2. b)
    METEOR (Banerjee and Lavie, 2005): Extends BLEU by incorporating synonym matching and recall:
    (10) $\text{METEOR}(A_{\text{LLM}},A_{\text{GT}})=F_{\text{mean}}\cdot(1-\text{Penalty})$
    where $F_{\text{mean}}$ balances precision and recall, and the penalty term accounts for word order discrepancies.
  3. c)
    ROUGE-L (Lin, 2004): Measures the longest common subsequence (LCS) overlap between $A_{\text{LLM}}$ and $A_{\text{GT}}$:
[…]
    where LCS denotes the longest matching word sequence.
@@ 4.1. Experimental Setup
Dataset Selection. We use the LAMD (Qian et al., 2025) dataset [^4] https://zenodo.org/records/14884736, which provides Android malware samples with high-quality GPT-4-generated ground-truth application purpose summaries, making it well-suited for evaluating semantic relevance. To reduce redundancy, we perform de-duplication by filtering out near-identical APKs within each malware category, resulting in 118 APKs across 6 categories (Adware, Backdoor, PUA, Riskware, Scareware, Trojan) and 13 families. All APKs are decompiled using Androguard, resulting in a total of $7,542,799$ decompiled functions across the dataset.
Implementation Details. For all selected models in Table 1, we use their official implementations from the Hugging Face Hub  [^5] https://huggingface.co/{meta-llama/CodeLlama-7b-Instruct-hf, HuggingFaceH4/starchat-beta, Salesforce/codet5-base-multi-sum, uclanlp/plbart-base} to ensure consistency and reproducibility. For models not originally instruction-tuned (i.e., CodeT5 and PLBART), we perform additional tuning on Java functions from their pretraining datasets. Each model is fine-tuned for 3 epochs, which we find sufficient to produce the structured outputs. The maliciousness score prediction of these two models is assisted by a locally deployed DeepSeek-R1-Distill-Llama-70B. In the fidelity evaluation, we use LightGBM (Ke et al., 2017) as the malware category classifier $C$, ensuring high reliability with an accuracy above $0.95$. To assess the effect of removing suspicious code, we experiment with top-$k$ values of $2$, $5$, and $8$. In the semantic relevance evaluation, to ensure stylistic consistency, we prompt all models to begin their outputs with the phrase: “This application appears to…”, matching the format used in the ground truth from LAMD. We set top-$v$ based on model context limits: $4$K tokens for CodeLlama, $8$K for StarChat, and $1$K for CodeT5 and PLBART. For BLEU-based evaluation, we use 2-gram precision, which is more appropriate for evaluating short summaries.
Our overall experimental results are summarized in subsubsection 3.3.3. In the following sections, we provide a detailed analysis of each research question: RQ1 in Section 4.2 and RQ2 in Section 4.3.
@@ 4.2. RQ1 - Decompiled Code
For maliciousness consistency, only CodeLlama and StarChat are evaluated, as CodeT5 and PLBART rely on external models for score generation. Among the two, StarChat achieves a notably higher score (over twice that of CodeLlama), suggesting a better understanding of high-level semantics relevant to malicious behavior. For name consistency, CodeLlama performs best, with StarChat following closely. Both models outperform the Seq2Seq baselines, reinforcing the observation that instruction-tuned, GPT-style models exhibit greater stability in structured output generation. Among the Seq2Seq models, PLBART outperforms CodeT5 by approximately $114\%$, likely due to its larger pretraining corpus and improved alignment between code and natural language.
In downstream evaluations, we observe clear performance differences across models in both fidelity and semantic relevance. StarChat consistently outperforms the others, demonstrating a stronger ability to assign meaningful maliciousness scores and produce high-level malware descriptions—results that align with its superior consistency metrics. This superior performance is likely driven by its larger model size and the StarCoder architecture, which emphasizes multilingual understanding and instruction-following, enhancing its ability to reason across diverse and obfuscated functions. CodeLlama performs competitively, especially excelling in top-2 function removal and producing stylistically aligned summaries, suggesting it effectively captures the most critical functions but is less robust than StarChat when evaluating broader function sets.
Among the Seq2Seq models, PLBART shows better performance in semantic relevance, benefiting from its BART-based architecture which favors fluent and coherent natural language generation. However, PLBART notably underperforms CodeT5 in fidelity, while both models’ maliciousness scores are generated externally by the same larger models. This difference arises because CodeT5’s summaries, though less fluent, contain more descriptions that better highlight critical code features, allowing the external scoring model to produce more discriminative maliciousness scores. Nevertheless, both Seq2Seq models exhibit limited stylistic control. For instance, even when explicitly prompted to begin with “This application appears to…”, they frequently prepend generic phrases like “This function…”, revealing limited control over stylistic constraints.
These findings highlight a fundamental tradeoff between linguistic fluency and semantic precision in LLM-generated outputs: for Seq2Seq models, while PLBART can produce readable summaries, they may not align well with underlying malicious behaviors; conversely, CodeT5 may better capture function-level semantics but lack expressive output formatting. Ultimately, our results reinforce the superiority of instruction-tuned GPT-style models (especially StarChat) for generating both accurate and interpretable outputs in fine-grained malware analysis tasks.
@@ 4.3. RQ2 - Function Naming
Overall, these findings indicate that LLM-based function renaming effectively enhances function-wise consistency and fidelity metrics, but may require careful handling to avoid diluting high-level semantic clarity. To mitigate this issue, future improvements could focus on calibrating the scores to better reflect model confidence and explicitly encoding more robust knowledge about malware semantics. Such enhancements may enable the models to maintain semantic clarity while benefiting from the improved consistency and fidelity introduced by meaningful function names.
@@ 5. Discussion
Cama opens the door to broader research directions. For example, it can support studies on malware concept drift (Pendlebury et al., 2019), enabling evaluation of whether Code LLM-based analyses generalize to evolving threats and unseen malware families. Beyond Code LLMs and Android malware, the framework is adaptable to assess a wide range of approaches, as long as they target core sub-tasks including function summarization, naming, and maliciousness estimation. Beyond benchmarking, Cama also supports practical applications: it can guide the selection, pretraining, or fine-tuning of Code LLMs specifically for malware tasks (Al-Kaswan et al., 2023; Yang et al., 2024; Shao et al., 2025). Its structured outputs, particularly maliciousness scores, can be used to prioritize suspicious functions, improving the precision of traditional malware classifiers (Arp et al., 2014; He et al., 2023a).
@@ 6. Conclusion
This paper introduces Cama, a benchmarking framework for systematically evaluating the effectiveness of open-source Code LLMs in Android malware analysis. We define a structured output format aligned with two key analysis tasks: malicious function identification and malware purpose summarization. To address the lack of fine-grained ground truth, we propose three domain-specific evaluation metrics, enabling rigorous assessment of LLM-generated outputs. Our benchmarking results reveal both the potential and current limitations of Code LLMs, emphasizing the necessity of structured evaluation frameworks to ensure stability and interpretability in real-world malware analysis. Cama provides a foundation for future work to select and adapt Code LLMs for malware analysis, improving their effectiveness in downstream tasks such as family classification and behavioral explanation.

### v2

@@ Abstract.
Large Language Models (LLMs) have demonstrated strong capabilities in various code intelligence tasks. However, their effectiveness for Android malware analysis remains underexplored. Decompiled Android malware code presents unique challenges for analysis, due to the malicious logic being buried within a large number of functions and the frequent lack of meaningful function names.
This paper presents Cama, a benchmarking framework designed to systematically evaluate the effectiveness of Code LLMs in Android malware analysis. Cama specifies structured model outputs to support key malware analysis tasks, including malicious function identification and malware purpose summarization. Built on these, it integrates three domain-specific evaluation metrics—consistency, fidelity, and semantic relevance—enabling rigorous stability and effectiveness assessment and cross-model comparison.
We construct a benchmark dataset of $118$ Android malware samples from $13$ families collected in recent years, encompassing over $7.5$ million distinct functions, and use Cama to evaluate four popular open-source Code LLMs. Our experiments provide insights into how Code LLMs interpret decompiled code and quantify the sensitivity to function renaming, highlighting both their potential and current limitations in malware analysis.
Code LLM, Malware Analysis
journalyear: 2025copyright: ccconference: 34th ACM SIGSOFT International Symposium on Software Testing and Analysis; June 25–28, 2025; Trondheim, Norwaybooktitle: 34th ACM SIGSOFT International Symposium on Software Testing and Analysis (ISSTA Companion ’25), June 25–28, 2025, Trondheim, Norwaydoi: 10.1145/3713081.3731745isbn: 979-8-4007-1474-0/2025/06ccs: Security and privacy Software and application security
@@ 1. Introduction
Recent advances in Large Language Models (LLMs) have transformed natural language processing (Radford et al., 2018), and their extension to code understanding has led to the emergence of Code LLMs, i.e., specialized LLMs trained on large-scale code repositories. These models have achieved strong performance in tasks such as code generation, summarization, and repair (Li et al., 2023b; Tian et al., 2024; Fang et al., 2024). Given their growing capabilities in reasoning about code, Code LLMs offer a promising direction for automating Android malware analysis, a domain where analysts must often manually inspect large volumes of low-level code to uncover malicious behavior (Mantovani et al., 2022).
While the potential is clear, effectively applying Code LLMs in this context remains challenging. First, decompiled Android code is often obfuscated, lacks type information, and contains incomplete control structures (Meng and Miller, 2016; Shang et al., 2024), diverging sharply from the clean, structured code these models are typically trained on (Zheng et al., 2023). Second, accurate semantic interpretation is difficult due to the high-level and diverse malicious behaviors in malware (Wang et al., 2022), while the absence of reliable function-level ground truth (e.g., labeling functions as malicious or benign) further complicates evaluation (He et al., 2023b; Kantchelian et al., 2015; Salem, 2021). These challenges underscore the need for a structured evaluation framework that systematically assesses and compares Code LLM performance in real-world malware analysis scenarios.
@@
For the first question, we analyze the quality of LLM-generated outputs using our structured output format and domain-specific metrics. For the second, we systematically rename functions in the original decompiled code with the LLM-suggested names and measure the subsequent impact on evaluation metrics. Our findings reveal that while Code LLMs can generate informative summaries, their understanding of maliciousness remains limited. Among existing Code LLMs, instruction-tuned GPT-style models significantly outperform Seq2Seq models across different metrics; function renaming enhances fidelity and consistency but may reduce semantic clarity, indicating a trade-off that warrants careful consideration.
@@ 2.1. Code Large Language Models
2.2. Learning-based Malware Analysis
LLM-powered Analysis. Recent efforts have explored LLM-powered malware detection, primarily operating within the established pipeline of conventional classifiers and leveraging LLMs in two ways: 1) querying LLMs with original features to generate detection outputs (Li et al., 2024), and 2) using LLMs to encode text-based semantic representations that enrich traditional feature spaces (Zhao et al., 2025). More advanced approaches leverage GPT-4o-mini’s code summarization capabilities (Walton et al., 2025), improving malware detection via program slicing techniques and multi-tiered factual checking (Qian et al., 2025). Despite these advancements, the effectiveness of Code LLMs in fine-grained analysis remains uncharted, largely due to the lack of ground truth. Our work goes beyond classification and complements existing studies by systematically benchmarking open-source Code LLMs on both function-level and app-level malware analysis tasks.
@@ 3.3.1. Consistency-based Metric
(1) $\text{JSD}(M^{\prime}_{\text{raw}},M^{\prime}_{\text{des}})=\frac{1}{2}D_{\text{KL}}(M^{\prime}_{\text{raw}}||M_{\text{avg}})+\frac{1}{2}D_{\text{KL}}(M^{\prime}_{\text{des}}||M_{\text{avg}})\,,$
@@ 3.3.3. Semantic-based Metric
We compare $A_{\text{LLM}}$ against the reference malware description $A_{\text{GT}}$ using three widely used text similarity metrics. BLEU (Papineni et al., 2002) measures n-gram precision between $A_{\text{LLM}}$ and $A_{\text{GT}}$:
@@
where $p_{n}$ is n-gram precision and $w_{n}$ are weighting factors. Additionally, we report METEOR (Banerjee and Lavie, 2005), which incorporates synonym matching and recall, and ROUGE-L (Lin, 2004), which evaluates the longest common subsequence overlap.
@@ 4.1. Experimental Setup
Dataset Selection. We use the LAMD (Qian et al., 2025) dataset [^4] LAMD dataset collected in the 2020s: https://zenodo.org/records/14884736, which provides Android malware samples with high-quality GPT-4-generated ground-truth application purpose summaries, making it well-suited for evaluating semantic relevance. To reduce redundancy, we perform de-duplication by filtering out near-identical APKs within each malware category, resulting in 118 APKs across 6 categories (Adware, Backdoor, PUA, Riskware, Scareware, Trojan) and 13 families. All APKs are decompiled using Androguard, resulting in a total of $7,542,799$ decompiled functions across the dataset.
Implementation Details. For all selected models in Table 1, we use their official implementations from the Hugging Face Hub  [^5] https://huggingface.co/{meta-llama/CodeLlama-7b-Instruct-hf, HuggingFaceH4/starchat-beta, Salesforce/codet5-base-multi-sum, uclanlp/plbart-base}. For models not originally instruction-tuned (i.e., CodeT5 and PLBART), we perform additional tuning on Java functions from their pretraining datasets. Each model is fine-tuned for 3 epochs, which we find sufficient to produce the structured outputs. The maliciousness score prediction of these two models is assisted by a locally deployed DeepSeek-R1-Distill-Llama-70B. In the fidelity evaluation, we use LightGBM (Ke et al., 2017) as the malware category classifier $C$, ensuring high reliability with an accuracy above $0.95$. To assess the effect of removing suspicious code, we experiment with top-$k$ values of $2$, $5$, and $8$. In the semantic relevance evaluation, to ensure stylistic consistency, we prompt all models to begin their outputs with the phrase: “This application appears to…”, matching the format used in the ground truth. We set top-$v$ based on model context limits: $4$K tokens for CodeLlama, $8$K for StarChat, and $1$K for CodeT5 and PLBART. For BLEU-based evaluation, we use 2-gram precision, which is more appropriate for evaluating short summaries.
The generated outputs are made open-source  [^6] Our Cama dataset: https://zenodo.org/records/15155917. Our overall experimental results are summarized in subsubsection 3.3.3. In the following sections, we provide a detailed analysis of each research question: RQ1 in Section 4.2 and RQ2 in Section 4.3.
@@ 4.2. RQ1 - Decompiled Code
For maliciousness consistency, only CodeLlama and StarChat are evaluated, as CodeT5 and PLBART rely on external models for score generation. Among the two, StarChat achieves a notably higher score (over twice that of CodeLlama), suggesting a better understanding of high-level malware semantics. For name consistency, CodeLlama performs best, with StarChat following closely. Both models outperform the Seq2Seq baselines, reinforcing the observation that instruction-tuned, GPT-style models exhibit greater stability in structured output generation. Among the Seq2Seq models, PLBART outperforms CodeT5 by approximately $114\%$, likely due to its improved alignment between code and natural language.
In downstream evaluations, we observe clear performance differences across models in both fidelity and semantic relevance. StarChat consistently outperforms the others, demonstrating a stronger ability to assign meaningful maliciousness scores and produce high-level malware descriptions—results that align with its superior consistency metrics. This superior performance is likely driven by its larger model size and the StarCoder architecture, which emphasizes multilingual understanding and instruction-following. CodeLlama performs competitively, especially excelling in top-2 function removal and producing stylistically aligned summaries, suggesting it effectively captures the most critical functions but is less robust than StarChat when evaluating broader function sets.
Among the Seq2Seq models, PLBART shows better performance in semantic relevance, benefiting from its BART-based architecture, which favors fluent and coherent natural language generation. However, PLBART notably underperforms CodeT5 in fidelity, while both models’ maliciousness scores are generated externally by the same larger models. This difference arises because CodeT5’s summaries, though less fluent, contain more descriptions that better highlight critical code features, allowing the external scoring model to produce more discriminative maliciousness scores. Nevertheless, both Seq2Seq models exhibit limited stylistic control. For instance, even when explicitly prompted to begin with “This application appears to…”, they frequently prepend generic phrases like “This function…”, revealing limited control over stylistic constraints.
These findings highlight a fundamental tradeoff between linguistic fluency and semantic precision in LLM-generated outputs. Our results also reinforce the superiority of instruction-tuned GPT-style models (especially StarChat) for generating both accurate and interpretable outputs in fine-grained malware analysis tasks.
@@ 4.3. RQ2 - Function Naming
Overall, these findings indicate that LLM-based function renaming effectively enhances function-wise consistency and fidelity metrics, but may require careful handling to avoid diluting high-level semantic clarity. To mitigate this issue, future improvements could focus on calibrating the scores to better reflect model confidence and explicitly encoding more robust knowledge about malware semantics.
@@ 5. Discussion
Cama opens the door to broader research directions. For example, it can support studies on malware concept drift (Pendlebury et al., 2019), enabling evaluation of whether Code LLM-based analyses generalize to evolving threats. Beyond Code LLMs and Android malware, the framework is adaptable to assess a wide range of approaches, as long as they target core sub-tasks including function summarization, naming, and maliciousness estimation. Beyond benchmarking, Cama also supports practical applications: it can guide the selection, pretraining, or fine-tuning of Code LLMs specifically for malware tasks (Al-Kaswan et al., 2023; Yang et al., 2024; Shao et al., 2025). Its structured outputs, particularly maliciousness scores, can be used to prioritize suspicious functions, improving the precision of traditional malware classifiers (Arp et al., 2014; He et al., 2023a).
@@ 6. Conclusion
This paper introduces Cama, a benchmarking framework for systematically evaluating the effectiveness of open-source Code LLMs in Android malware analysis. We define a structured output format aligned with two key analysis tasks: malicious function identification and malware purpose summarization. To address the lack of fine-grained ground truth, we propose three domain-specific evaluation metrics, enabling rigorous assessment of LLM-generated outputs. Our benchmarking results reveal both the potential and current limitations of Code LLMs. Cama provides a foundation for future work to select and adapt Code LLMs for malware analysis, improving their effectiveness in downstream tasks such as family classification and behavioral explanation.